# CellCLIP – Learning Perturbation Effects in Cell Painting via Text-Guided Contrastive Learning

### v2

@@ Abstract
High-content screening (HCS) assays based on high-throughput microscopy techniques such as Cell Painting have enabled the interrogation of cells’ morphological responses to perturbations at an unprecedented scale. The collection of such data promises to facilitate a better understanding of the relationships between different perturbations and their effects on cellular state. Towards achieving this goal, recent advances in cross-modal contrastive learning could, in theory, be leveraged to learn a unified latent space that aligns perturbations with their corresponding morphological effects. However, the application of such methods to HCS data is not straightforward due to substantial differences in the semantics of Cell Painting images compared to natural images, and the difficulty of representing different classes of perturbations (e.g. small molecule vs CRISPR gene knockout) in a single latent space. In response to these challenges, here we introduce CellCLIP, a cross-modal contrastive learning framework for HCS data. CellCLIP leverages pre-trained image encoders coupled with a novel channel encoding scheme to better capture relationships between different microscopy channels in image embeddings, along with natural language encoders for representing perturbations. Our framework outperforms current open-source models, demonstrating the best performance in both cross-modal retrieval and biologically meaningful downstream tasks while also achieving significant reductions in computation time. Code for our reproducing our experiments is available at https://github.com/suinleelab/CellCLIP/tree/main.
@@ 1 Introduction
A grand challenge in cellular biology is understanding the impacts of different perturbations, such as exposure to chemical compounds or gene knockouts, on cellular function. In pursuit of this goal, a number of high-content screening (HCS) techniques have been developed that combine image-based deep phenotyping with high-throughput perturbations [20, 29]. For example, the Cell Painting protocol [4] stains cells with six fluorescent dyes highlighting distinct cellular components (e.g. Hoescht staining for nuclear DNA), which are then imaged in five channels via automated microscopy (Figure 1a). Cell Painting is now routinely combined with other biotechnological advances to measure cells’ response to thousands of chemical and/or genetic perturbations in parallel at low cost [39, 35].
@@
To address these issues, here we propose CellCLIP (Figure 1b), a CL framework designed for the unique challenges in Cell Painting perturbation data. Our framework employs off-the-shelf pretrained vision models alongside a novel channel-aware encoding scheme (Figure 1c) to facilitate reasoning across the different sources of information captured in Cell Painting channels; in addition, our image encoding architecture incorporates techniques from the multiple-instance-learning literature [24] to account for the many-to-one nature of Cell Painting perturbation-image pairs. To represent images’ corresponding perturbations, CellCLIP employs natural language encoders, enabling application to multiple classes of perturbations with a single model. We benchmarked CellCLIP on profile-perturbation retrieval for unseen compounds [4] and its ability to recover known biological relationships between perturbations [10, 27]. We find that CellCLIP achieves strong performance compared to the previous state-of-the-art while taking a fraction of the time to train. Our method also shows promising results in cross-perturbation class matching, thereby providing a scalable and effective solution for biological discovery.
@@ 2 Background
$\mathcal{L}_{\mathrm{CLIP}}=\frac{1}{N}\sum_{i=1}^{N}\Bigg{[}\underbrace{-\log\frac{\exp(\langle p_{i}\cdot q_{i}\rangle/\tau)}{\sum_{j=1}^{N}\exp(\langle p_{i}\cdot q_{j}\rangle/\tau)}}_{\mathcal{U}\rightarrow\mathcal{V}}+\underbrace{-\log\frac{\exp(\langle p_{i}\cdot q_{i}\rangle/\tau)}{\sum_{j=1}^{N}\exp(\langle p_{j}\cdot q_{i}\rangle/\tau)}}_{\mathcal{V}\rightarrow\mathcal{U}}\Bigg{]}$ (1)
@@ 3 Methods
Here we present CellCLIP in detail. We first present CrossChannelFormer, our image encoding scheme that facilitates CL with Cell Painting data (Section 3.1); due to the idiosyncracies of Cell Painting perturbation datasets, this encoding must be done with care, and we cannot naively reuse strategies for CL with natural images. We then present our approach for encoding perturbation labels (Section 3.2) and outline our training strategy (Section 3.3).
@@ 3.1 CrossChannelFormer
Our CrossChannelFormer architecture for image encoding (Figure 1c) consists of three steps that allow us to effectively account for the nuances of Cell Painting perturbation data without introducing prohibitive computational costs. We describe these steps below.
@@
that we employ as an image’s profile. In our experiments, we used a frozen, pretrained DINOv2 (giant) model for our feature extractor $\psi$, and explore the impact of different choices for $\psi$ in Section A.1.
@@
where $\mathcal{S}(\cdot)$ denotes a permutation-invariant transformation function applied channel-wise (e.g., a mean computed separately for each channel). For our experiments we use the gated attention pooling operator of Ilse et al. [24] for $\mathcal{S}$, and we explore the impact of different choices for $\mathcal{S}$ in Section A.2. In addition to improving performance on downstream tasks, per-perturbation pooling also improves computational efficiency: instead of computing a contrastive loss for each of the $N_{i}$ individual profiles associated with a perturbation $i$, we compute a single pooled representation. This reduces the number of positive pairs per perturbation from $N_{i}$ to one, leading to an approximate $N_{i}$-fold speedup in training time compared to instance-level supervision.
@@
Thus, to enable cross-channel reasoning while minimizing computational costs, we introduce CrossChannelFormer, a specialized encoder architecture for CellCLIP. Unlike the standard Vision Transformer (ViT; [1]), where each input token represents a multi-channel image patch, CrossChannelFormer takes pooled profiles $\mu(U_{i})$ (Equation 3) that capture the global cellular features associated with specific stains. We then introduce a set of learnable channel embeddings, $[\texttt{chn}^{1},\dots,\texttt{chn}^{C}]$, where each $\texttt{chn}^{c}\in\mathbb{R}^{d}$ encodes information unique to its respective channel. Finally, we prepend a learnable classifier token $\texttt{cls}\in\mathbb{R}^{d}$ to the sequence, which aggregates global image features across all channels. The resulting input sequence to the transformer is:
$[\texttt{cls},\mu(U_{i})^{1}+\texttt{chn}^{1},\mu(U_{i})^{2}+\texttt{chn}^{2},\dots,\mu(U_{i})^{C}+\texttt{chn}^{C}],$ (4)
@@ 3.2 Perturbation Encoding via Natural Language
For our experiments, we use a pretrained BERT model [12] which uses the WordPiece tokenizer [44] and supports 512 tokens. The prompt template used is provided in Appendix B.
@@ 3.3 CellCLIP Training Objective
In standard contrastive learning, the objective is to align matched data pairs by minimizing the loss in Equation 1. In Cell Painting, in addition to cross-modal retrieval [16, 37], many intra-modal downstream tasks, such as matching related perturbations [10, 25, 27], are also important. However, maintaining adequate performance on intra-modal tasks is non-trivial; for example, in the context of Cell Painting perturbation screens, many distinct perturbations induce highly similar morphological profiles, i.e., $f_{\theta^{*}}(\mu(U_{i}))\approx f_{\theta^{*}}(\mu(U_{j}))$ for $i\neq j$, where $f_{\theta^{*}}$ is an oracle profile encoder. In this setting, treating all non-matching pairs as hard negatives can degrade retrieval performance by encouraging too much separation between biologically similar perturbations.
@@
To address this, we adopt the Continuously Weighted Contrastive Loss (CWCL) [40] in CellCLIP. CWCL replaces binary labels with continuous similarity-based weights (Figure 1b), applying a soft labeling scheme within the cross-entropy loss. The loss helps training cross-modal models by ensuring that similarities picked up by a strong pretrained unimodal embedding model are maintained in the cross-modal representation space. In our case, we apply the CWCL loss by computing target similarities between per-perturbation pooled profiles so that morphologically similar profiles remain close in CellCLIP’s embedding space. For the profile-to-perturbation direction, $\mathcal{U}\rightarrow\mathcal{V}$, our adapted CWCL objective is given by
@@
$\mathcal{L}_{\mathrm{CWCL},\mathcal{U}\rightarrow\mathcal{V}}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{\sum_{j\in[N]}w_{ij}^{\mathcal{U}}}\left[\sum_{j=1}^{N}w_{ij}^{\mathcal{U}}\cdot\log\frac{\exp(\langle p_{i}\cdot q_{i}\rangle/\tau)}{\sum_{k=1}^{N}\exp(\langle p_{i}\cdot q_{k}\rangle/\tau)}\right].$ (5)
@@
Here $p_{i}$ represents the output of CrossChannelFormer applied to images corresponding to perturbation $i$, $q_{i}$ corresponds to our encoding of the natural language description of perturbation $i$, and $w_{ij}^{\mathcal{U}}$ denotes the similarity between the pooled profiles $\mu(U_{i})$ and $\mu(U_{j})$, used to reweight the alignment from modality $\mathcal{U}$ to modality $\mathcal{V}$. We compute $w_{ij}^{\mathcal{U}}$ as the average channel-wise cosine similarity between pooled profiles. [^2] Specifically, $w_{ij}^{\mathcal{U}}=\sum_{c=1}^{C}\frac{\langle\mu(U_{i})^{c},\mu(U_{j})^{c}\rangle}{2C}+0.5$, ensuring $w_{ij}^{\mathcal{U}}\in[0,1]$. In Section A.3 we provide results illustrating the benefits of our choice over other potential contrastive losses.
@@
For the perturbation-to-profile direction, $\mathcal{V}\rightarrow\mathcal{U}$, we apply the standard CLIP loss (Equation 1) as we found computing similarities either within the input space $\mathcal{V}$ (i.e., discrete tokens) or in the projected space $\mathcal{Q}$ (i.e., BERT embeddings) did not provide meaningful signals for soft-label supervision. Thus, the final training objective is the sum of the two directional losses:
@@ 4.2 Intra-Modality Retrieval on Biologically Meaningful Tasks
Retrieval performance is evaluated against ground-truth labels via task-specific metrics used in previous work. For replicate detection and sister perturbation matching, we report mean average precision (mAP) as in Chandrasekaran et al. [10]. For biological relationship recovery, we follow Kraus et al. [27] and report recall over the top and bottom 5% of the similarity distribution (i.e., 10% recall). Further details regarding the computation of these metrics are provided in Appendix C.
@@ 4.3 Datasets
For cross-modality retrieval, following Sanchez-Fernandez et al. [37], we utilize a curated version of data from Bray et al. [4]. This dataset has 759,782 five-channel Cell Painting images corresponding to 30,404 small-molecule perturbations. We partitioned the dataset by perturbation into train, validation, and test sets with a 70/10/20 split, resulting in 2115 unseen small molecules in the test set.
For replicate detection & sister perturbation matching, we employ CPJUMP1 [10], which features 186,925 nine-channel microscopy images.These images include three bright-field channels in addition to six Cell Painting dye channels. They are perturbed across 650 distinct perturbations, including compounds and genetic modifications such as CRISPR and ORF interventions. For each perturbation class, we applied a 70/10/20 split for training, validation, and testing.
For zero-shot gene-gene relationship recovery, we use RxRx3-core [27], a curated subset of RxRx3 [15], covering 736 gene knockouts. Gene-gene relationships are obtained from a set of public databases (Reactome, HuMAP, SIGNOR, StringDB, and CORUM; Table S.7) as in Kraus et al. [26].
Further details on datasets and preprocessing are provided in Appendix D.
@@ 4.4 Model training
For cross-modality retrieval evaluation on Bray et al. [4], CellCLIP was trained with 50 epochs using a batch size of 512 and an AdamW optimizer. The learning rate was set at $2\times 10^{-4}$ with cosine annealing and restart. The temperature parameter, $\tau$, is initialized to 14.3. For replicate detection and sister perturbation matching in CPJUMP1 [10], we reused our model trained on the Bray et al. [4] dataset and fine-tuned it for another 50 epochs, using the same parameter settings as during their initial training. For gene relationship recovery using RxRx3-core, we perform this evaluation in a “zero-shot” manner and evaluate models trained on Bray et al. [4] without subsequently fine-tuning on RxRx3-core. More details about training CellCLIP and other baselines can be found in Appendix B.
@@
[…]
@@ 5.1 Cross-modality retrieval
We present our results for cross-modality retrieval tasks (i.e., perturbation-to-profile and profile-to-perturbation) in Table 1. For our benchmarking we compared CellCLIP against previous state-of-the-art CL methods for Cell Painting perturbation screens: CLOOME [37] and MolPhenix [16]. For CLOOME we used the same architecture with optimal hyperparameters as specified in Sanchez-Fernandez et al. [37], along with subsequent variations of the model described in Fradkin et al. [16]. Because Fradkin et al. [16]’s implementation of MolPhenix relied on closed-source components trained on non-public datasets, compared against an approximation of their framework which we designate as MolPhenix*. We refer to Appendix E for further details.
Overall, we found that CellCLIP demonstrated substantially higher performance at both cross-modal retrieval tasks compared to baseline methods. To understand the source of performance gains in CellCLIP, we conducted a series of ablations to understand the contribution of each component in CellCLIP to retrieval performance. Specifically, starting with CLOOME’s proposed encoding scheme (Appendix E), where individual images encoded using ResNet50 are aligned with chemical perturbations encoded using Morgan fingerprints combined with an MLP, we gradually replaced each of CLOOME’s components with those of CellCLIP and assessed each change’s impact on model performance (Table 2). We describe our findings from these experiments in detail below.
@@
We next investigated the impact of CrossChannelFormer’s ability to reason across global Cell Painting channel information. To do so, we replaced CLOOME’s ResNet image encoder with our CrossChannelFormer encoder. To isolate the effects of CrossChannelFormer’s ability to reason across channels from the effects of per-perturbation pooling, in this experiment we removed the CrossChannelFormer’s pooling function and trained the resulting model (denoted as CrossChannelFormer in Table 2) on individual image-perturbation pairs.
@@
Finally, we evaluated the impact of the attention-based pooling operator [24] within CrossChannelFormer compared to instance-level training (i.e., no pooling). We found that including pooling resulted in yet another increase in model performance. In addition, by reducing the number of pairs for contrastive loss computation, pooling yielding a 6.7 times speedup in training time relative to instance-level training. We also explored the impact of varying our choice of pooling operator (Section A.2), and found that attention-based pooling yielded the best results among the pooling operators we considered.

### v3

@@ Abstract
High-content screening (HCS) assays based on high-throughput microscopy techniques such as Cell Painting have enabled the interrogation of cells’ morphological responses to perturbations at an unprecedented scale. The collection of such data promises to facilitate a better understanding of the relationships between different perturbations and their effects on cellular state. Towards achieving this goal, recent advances in cross-modal contrastive learning could, in theory, be leveraged to learn a unified latent space that aligns perturbations with their corresponding morphological effects. However, the application of such methods to HCS data is not straightforward due to substantial differences in the semantics of Cell Painting images compared to natural images, and the difficulty of representing different classes of perturbations (e.g. small molecule vs CRISPR gene knockout) in a single latent space. In response to these challenges, here we introduce CellCLIP, a cross-modal contrastive learning framework for HCS data. CellCLIP leverages pre-trained image encoders coupled with a novel channel encoding scheme to better capture relationships between different microscopy channels in image embeddings, along with natural language encoders for representing perturbations. Our framework outperforms current open-source models, demonstrating the best performance in both cross-modal retrieval and biologically meaningful downstream tasks while also achieving significant reductions in computation time. Code for our reproducing our experiments is available at https://github.com/suinleelab/CellCLIP.
@@ 1 Introduction
A grand challenge in cellular biology is understanding the impacts of different perturbations, such as exposure to chemical compounds or gene knockouts, on cellular function. In pursuit of this goal, a number of high-content screening (HCS) techniques have been developed that combine image-based deep phenotyping with high-throughput perturbations [20, 30]. For example, the Cell Painting protocol [4] stains cells with six fluorescent dyes highlighting distinct cellular components (e.g. Hoescht staining for nuclear DNA), which are then imaged in five channels via automated microscopy (Figure˜1a). Cell Painting is now routinely combined with other biotechnological advances to measure cells’ response to thousands of chemical and/or genetic perturbations in parallel at low cost [40, 36].
@@
To address these issues, here we propose CellCLIP (Figure˜1b), a CL framework designed for the unique challenges in Cell Painting perturbation data. Our framework employs off-the-shelf pretrained vision models alongside a novel channel-aware encoding scheme (Figure˜1c) to facilitate reasoning across the different sources of information captured in Cell Painting channels; in addition, our image encoding architecture incorporates techniques from the multiple-instance-learning literature [24] to account for the many-to-one nature of Cell Painting perturbation-image pairs. To represent images’ corresponding perturbations, CellCLIP employs natural language encoders, enabling application to multiple classes of perturbations with a single model. We benchmarked CellCLIP on profile-perturbation retrieval for unseen compounds [4] and its ability to recover known biological relationships between perturbations [10, 28]. We find that CellCLIP achieves strong performance compared to the previous state-of-the-art while taking a fraction of the time to train. Our method also shows promising results in cross-perturbation class matching, thereby providing a scalable and effective solution for biological discovery.
@@ 2 Background
$\mathcal{L}_{\mathrm{CLIP}}=\frac{1}{N}\sum_{i=1}^{N}\Bigg[\underbrace{-\log\frac{\exp(\langle p_{i}\cdot q_{i}\rangle/\penalty 50\tau)}{\sum_{j=1}^{N}\exp(\langle p_{i}\cdot q_{j}\rangle/\penalty 50\tau)}}_{\mathcal{U}\rightarrow\mathcal{V}}+\underbrace{-\log\frac{\exp(\langle p_{i}\cdot q_{i}\rangle/\penalty 50\tau)}{\sum_{j=1}^{N}\exp(\langle p_{j}\cdot q_{i}\rangle/\penalty 50\tau)}}_{\mathcal{V}\rightarrow\mathcal{U}}\Bigg]$ (1)
@@ 3 Methods
Here we present CellCLIP in detail. We first present CrossChannelFormer, our image encoding scheme that facilitates CL with Cell Painting data (Section˜3.1); due to the idiosyncracies of Cell Painting perturbation datasets, this encoding must be done with care, and we cannot naively reuse strategies for CL with natural images. We then present our approach for encoding perturbation labels (Section˜3.2) and outline our training strategy (Section˜3.3).
@@ 3.1 CrossChannelFormer
Our CrossChannelFormer architecture for image encoding (Figure˜1c) consists of three steps that allow us to effectively account for the nuances of Cell Painting perturbation data without introducing prohibitive computational costs. We describe these steps below.
@@
that we employ as an image’s profile. In our experiments, we used a frozen, pretrained DINOv2 (giant) model for our feature extractor $\psi$, and explore the impact of different choices for $\psi$ in Section˜A.1.
@@
where $\mathcal{S}(\cdot)$ denotes a permutation-invariant transformation function applied channel-wise (e.g., a mean computed separately for each channel). For our experiments we use the gated attention pooling operator of Ilse et al. [24] for $\mathcal{S}$, and we explore the impact of different choices for $\mathcal{S}$ in Section˜A.2. In addition to improving performance on downstream tasks, per-perturbation pooling also improves computational efficiency: instead of computing a contrastive loss for each of the $N_{i}$ individual profiles associated with a perturbation $i$, we compute a single pooled representation. This reduces the number of positive pairs per perturbation from $N_{i}$ to one, leading to an approximate $N_{i}$-fold speedup in training time compared to instance-level supervision.
@@
Thus, to enable cross-channel reasoning while minimizing computational costs, we introduce CrossChannelFormer, a specialized encoder architecture for CellCLIP. Unlike the standard Vision Transformer (ViT; [1]), where each input token represents a multi-channel image patch, CrossChannelFormer takes pooled profiles $\mu(U_{i})$ (Equation˜3) that capture the global cellular features associated with specific stains. We then introduce a set of learnable channel embeddings, $\texttt{chn}^{1},\dots,\texttt{chn}^{C}]$, where each $\texttt{chn}^{c}\in\mathbb{R}^{d}$ encodes information unique to its respective channel. Finally, we prepend a learnable classifier token $\texttt{cls}\in\mathbb{R}^{d}$ to the sequence, which aggregates global image features across all channels. The resulting input sequence to the transformer is:
$\texttt{cls},\mu(U_{i})^{1}+\texttt{chn}^{1},\mu(U_{i})^{2}+\texttt{chn}^{2},\dots,\mu(U_{i})^{C}+\texttt{chn}^{C}],$ (4)
@@ 3.2 Perturbation Encoding via Natural Language
For our experiments, we use a pretrained BERT model [12] which uses the WordPiece tokenizer [46] and supports 512 tokens. The prompt template used is provided in Appendix˜B.
@@ 3.3 CellCLIP Training Objective
In standard contrastive learning, the objective is to align matched data pairs by minimizing the loss in Equation˜1. In Cell Painting, in addition to cross-modal retrieval [16, 38], many intra-modal downstream tasks, such as matching related perturbations [10, 25, 28], are also important. However, maintaining adequate performance on intra-modal tasks is non-trivial; for example, in the context of Cell Painting perturbation screens, many distinct perturbations induce highly similar morphological profiles, i.e., $f_{\theta^{*}}(\mu(U_{i}))\approx f_{\theta^{*}}(\mu(U_{j}))$ for $i\neq j$, where $f_{\theta^{*}}$ is an oracle profile encoder. In this setting, treating all non-matching pairs as hard negatives can degrade retrieval performance by encouraging too much separation between biologically similar perturbations.
@@
To address this, we adopt the Continuously Weighted Contrastive Loss (CWCL) [41] in CellCLIP. CWCL replaces binary labels with continuous similarity-based weights (Figure˜1b), applying a soft labeling scheme within the cross-entropy loss. The loss helps training cross-modal models by ensuring that similarities picked up by a strong pretrained unimodal embedding model are maintained in the cross-modal representation space. In our case, we apply the CWCL loss by computing target similarities between per-perturbation pooled profiles so that morphologically similar profiles remain close in CellCLIP’s embedding space. For the profile-to-perturbation direction, $\mathcal{U}\rightarrow\mathcal{V}$, our adapted CWCL objective is given by
@@
$\mathcal{L}_{\mathrm{CWCL},\mathcal{U}\rightarrow\mathcal{V}}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{\sum_{j\in[N]}w_{ij}^{\mathcal{U}}}\left[\sum_{j=1}^{N}w_{ij}^{\mathcal{U}}\cdot\log\frac{\exp(\langle p_{i}\cdot q_{j}\rangle/\penalty 50\tau)}{\sum_{k=1}^{N}\exp(\langle p_{i}\cdot q_{k}\rangle/\penalty 50\tau)}\right].$ (5)
@@
Here $p_{i}$ represents the output of CrossChannelFormer applied to images corresponding to perturbation $i$, $q_{i}$ corresponds to our encoding of the natural language description of perturbation $i$, and $w_{ij}^{\mathcal{U}}$ denotes the similarity between the pooled profiles $\mu(U_{i})$ and $\mu(U_{j})$, used to reweight the alignment from modality $\mathcal{U}$ to modality $\mathcal{V}$. We compute $w_{ij}^{\mathcal{U}}$ as the average channel-wise cosine similarity between pooled profiles. [^2] Specifically, $w_{ij}^{\mathcal{U}}=\sum_{c=1}^{C}\frac{\langle\mu(U_{i})^{c},\mu(U_{j})^{c}\rangle}{2C}+0.5$, ensuring $w_{ij}^{\mathcal{U}}\in[0,1]$. In Section˜A.4 we provide results illustrating the benefits of our choice over other potential contrastive losses.
@@
For the perturbation-to-profile direction, $\mathcal{V}\rightarrow\mathcal{U}$, we apply the standard CLIP loss (Equation˜1) as we found computing similarities either within the input space $\mathcal{V}$ (i.e., discrete tokens) or in the projected space $\mathcal{Q}$ (i.e., BERT embeddings) did not provide meaningful signals for soft-label supervision. Thus, the final training objective is the sum of the two directional losses:
@@ 4.2 Intra-Modality Retrieval on Biologically Meaningful Tasks
Retrieval performance is evaluated against ground-truth labels via task-specific metrics used in previous work. For replicate detection and sister perturbation matching, we report mean average precision (mAP) as in Chandrasekaran et al. [10]. For biological relationship recovery, we follow Kraus et al. [28] and report recall over the top and bottom 5% of the similarity distribution (i.e., 10% recall). Further details regarding the computation of these metrics are provided in Appendix˜C.
@@ 4.3 Datasets
For cross-modality retrieval, following Sanchez-Fernandez et al. [38], we utilize a curated version of Bray et al. [4], comprising approximately 284,034 five-channel Cell Painting images corresponding to 10,578 small-molecule perturbations. We partitioned the dataset by perturbation into train, validation, and test sets with a 70/10/20 split, resulting in 2,115 unseen small molecules in the test set.
For replicate detection & sister perturbation matching, we employ CPJUMP1 [10], which features 186,925 eight-channel microscopy images. These images include three bright-field channels in addition to five Cell Painting dye channels. They are perturbed across 650 distinct perturbations, including compounds and genetic modifications such as CRISPR and ORF interventions. For each perturbation class, we applied a 70/10/20 split for training, validation, and testing.
For zero-shot gene-gene relationship recovery, we use RxRx3-core [28], a curated subset of RxRx3 [15], covering 736 gene knockouts. Gene-gene relationships are obtained from a set of public databases (Reactome, HuMAP, SIGNOR, StringDB, and CORUM; Table˜S.10) as in Kraus et al. [27].
Further details on datasets and preprocessing are provided in Appendix˜D.
@@ 4.4 Model training
For cross-modality retrieval evaluation on Bray et al. [4], CellCLIP was trained with 50 epochs using a batch size of 512 and an AdamW optimizer. The learning rate was set at $2\times 10^{-4}$ with cosine annealing and restart. The temperature parameter, $\tau$, is initialized to 14.3. For replicate detection and sister perturbation matching in CPJUMP1 [10], we reused our model trained on the Bray et al. [4] dataset and fine-tuned it for another 50 epochs, using the same parameter settings as during their initial training. For gene relationship recovery using RxRx3-core, we perform this evaluation in a “zero-shot” manner and evaluate models trained on Bray et al. [4] without subsequently fine-tuning on RxRx3-core. More details about training CellCLIP and other baselines can be found in Appendix˜B.
@@
$\ddagger$} Indicates that CLOOME uses the same image profile encoding procedure and architecture as MolPhenix*
@@ 5.1 Cross-modality retrieval
We present our results for cross-modality retrieval tasks (i.e., perturbation-to-profile and profile-to-perturbation) in Table˜1. For our benchmarking we compared CellCLIP against previous state-of-the-art CL methods for Cell Painting perturbation screens: CLOOME [38] and MolPhenix [16]. For CLOOME we used the same architecture with optimal hyperparameters as specified in Sanchez-Fernandez et al. [38], along with subsequent variations of the model described in Fradkin et al. [16]. Because Fradkin et al. [16]’s implementation of MolPhenix relied on closed-source components trained on non-public datasets, compared against an approximation of their framework which we designate as MolPhenix*. We refer to Appendix˜E for further details.
Overall, we found that CellCLIP demonstrated substantially higher performance at both cross-modal retrieval tasks compared to baseline methods. To understand the source of performance gains in CellCLIP, we conducted a series of ablations to understand the contribution of each component in CellCLIP to retrieval performance. Specifically, starting with CLOOME’s proposed encoding scheme (Appendix˜E), where individual images encoded using ResNet50 are aligned with chemical perturbations encoded using Morgan fingerprints combined with an MLP, we gradually replaced each of CLOOME’s components with those of CellCLIP and assessed each change’s impact on model performance (Table˜2). We describe our findings from these experiments in detail below.
@@
Table 3: Retrieval performance of CellCLIP trained with various perturbation-related prompts. Recall@1, @5, and @10 $\pm$ standard deviation across random seeds for perturbation-to-profile and profile-to-perturbation retrieval tasks.
Prompt modification Perturb-to-profile (%) Profile-to-perturb (%)
R@1 R@5 R@10 R@1 R@5 R@10
Original $1.18\pm 0.20$ $4.49\pm 0.06$ $7.37\pm 0.20$ $1.25\pm 0.10$ $4.82\pm 0.10$ $7.39\pm 0.23$
Without SMILES string $0.13\pm 0.05$ $0.68\pm 0.19$ $1.26\pm 0.14$ $0.14\pm 0.07$ $0.63\pm 0.17$ $1.24\pm 0.13$
Without cell type information $1.11\pm 0.13$ $4.02\pm 0.25$ $6.59\pm 0.33$ $1.10\pm 0.23$ $4.07\pm 0.29$ $6.36\pm 0.32$
Without drug name $1.47\pm 0.29$ $4.77\pm 0.20$ $7.37\pm 0.29$ $1.61\pm 0.26$ $4.89\pm 0.19$ $7.21\pm 0.27$
With DrugBank description $1.13\pm 0.16$ $4.26\pm 0.43$ $6.94\pm 0.37$ $1.16\pm 0.17$ $4.34\pm 0.36$ $6.71\pm 0.26$
Effect of perturbation prompts.
We next investigated alternative prompting strategies to assess the contribution of individual components. We re-trained CellCLIP after removing specific elements of the prompt template (Section˜3.2), such as cell type, drug name, SMILES string, one at a time (Table˜3). We found that removing the SMILES string caused a substantial drop in performance, while removing cell type information produced a smaller degradation. In contrast, removing drug names had little effect, likely because the richer SMILES representation already captures most of the relevant information.
We also tested whether prompts could be enriched with external metadata by incorporating DrugBank descriptions [26] for compounds present in the database. However, this did not yield performance improvements over the original template. A likely reason is the limited coverage of DrugBank, which includes only about 10% of the compounds in Bray et al. [4], as many of them are experimental molecules. Overall, these findings suggest that our original prompting strategy is already highly effective.
@@
We next investigated the impact of CrossChannelFormer’s ability to reason across global Cell Painting channel information. To do so, we replaced CLOOME’s ResNet image encoder with our CrossChannelFormer encoder. To isolate the effects of CrossChannelFormer’s ability to reason across channels from the effects of per-perturbation pooling, in this experiment we removed the CrossChannelFormer’s pooling function and trained the resulting model (denoted as CrossChannelFormer in Table˜2) on individual image-perturbation pairs.
@@
Finally, we evaluated the impact of the attention-based pooling operator [24] within CrossChannelFormer compared to instance-level training (i.e., no pooling). We found that including pooling resulted in yet another increase in model performance. In addition, by reducing the number of pairs for contrastive loss computation, pooling yielding a 6.7 times speedup in training time relative to instance-level training. We also explored the impact of varying our choice of pooling operator (Section˜A.2), and found that attention-based pooling yielded the best results among the pooling operators we considered. Moreover, to validate that the attention pooling learns to prioritize cells that most strongly reflect the perturbation effect, we performed a top-$n$ removal experiment: for each perturbation, we removed the $n$ samples with the highest attention weights. This led to a substantial performance drop (Table˜S.3), demonstrating that attention pooling effectively identifies and emphasizes the subset of cells carrying the strongest perturbation signal.